OpenStudy (anonymous):

how do you write that on a graph? y = − x2 + 3 y = x2 −3 fan and medal

OpenStudy (anonymous):

\[y = − x^2 + 3 \\ y = x^2 −3 \] two parabolas one open up one open down

OpenStudy (anonymous):

@satellite73 would they intersect?

OpenStudy (anonymous):

yes, the one that opens down \(y=-x^2+3\) has vertex \((0,0)\) and the one that opens up \(y=x^2-3\) has vertex at \((0,-3)\)

OpenStudy (midhun.madhu1987):

 Try to plot these points on the graph and you can check it..